If two buildings are 30m m apart then with what speed a ball must be thrown from a window which is at a height 150 m above the ground so that it reaches a window which is at a height of a27.5 m above the ground

Grade:11

1 Answers

Vikas TU
7 years ago
150 – 27.5 = 4.9t^2
122.5 = 4.9t^2
t^2 = 25
t =5 s
 
it will cover  30 m that is horizontal ditasnce and a= 0 for that
thus,
30 = u*5
u = 6 m/s